-2(-7k+4)+9=-13
I need to find k.

Answer by fractalier(6550)

Okay, from
-2(-7k+4)+9=-13
let's expand the left-hand side and get
14k - 8 + 9 = -13
Now combine
14k + 1 = -13
Now subtract 1, and get
14k = -14
and divide by 14 to get
k = -1